# The low dimensional homology groups of the elementary group of degree two

### v2

@@ Abstract.
In this article we study the first, the second and the third homology groups of the elementary group $\mathit{{\rm E}}_{2}(A)$, where $A$ is a commutative ring. In particular, we prove a refined Bloch-Wigner type exact sequence over a semilocal ring (with some mild restriction on its residue fields) such that $-1\in({A^{\times}})^{2}$ or $|{A^{\times}}/({A^{\times}})^{2}|\leq 4$.
For a commutative ring $A$ with unit, let $\mathit{{\rm E}}_{2}(A)$ be the subgroup of the general linear group $\mathit{{\rm GL}}_{2}(A)$ generated by elementary matrices. Clearly $\mathit{{\rm E}}_{2}(A)\subseteq\mathit{{\rm SL}}_{2}(A)$. We say that $A$ is a ${\rm GE}_{2}(A)$-ring if $\mathit{{\rm E}}_{2}(A)=\mathit{{\rm SL}}_{2}(A)$.
In two previous papers [16] and [17] we studied the low dimensional homology groups of the elementary group $\mathit{{\rm E}}_{2}(A)$. In this paper, we further study these homology groups.
The first main result of this paper is over the first homology group of $\mathit{{\rm E}}_{2}(A)$. We show that for any commutative ring $A$, we have the exact sequence
@@
where $Y_{\bullet}(A^{2})$ is a complex of ${\rm GE}_{2}(A)$-modules obtained from certain unimodular vectors in $A^{2}$ (see Theorem 4.1). Here $M$ is the additive subgroup of $A$ generated by $x(a^{2}-1)$ and $3(b+1)(c+1)$, where $x\in A$ and $a,b,c\in{A^{\times}}$. The right map is induced by the map $A\rightarrow\mathit{{\rm E}}_{2}(A)$, where sends $x$ to the elementary matrix $E_{12}(x)$. The group $A/M$ and its connection with the first homology of $\mathit{{\rm E}}_{2}(A)$ first were studied by Cohn in [3] and [4].
Our second main result concerns the second homology group $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$. It is easily to show that this groups sits in the exact sequence
@@
where $\mathrm{B}(A)$ is the subgroup of $\mathit{{\rm E}}_{2}(A)$ consisting of all matrices with lower left hand entry equal to zero and $I(A)$ is the fundamental ideal of $A$. Our second result concerns the map $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)$ appeared in the above exact sequence. When $2\in{A^{\times}}$ we find cycle that their homology class in $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ maps to the elements $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle$, $a,b\in{A^{\times}}$ (see Proposition 5.2 for the precise statement).
The third main result of this paper is a refined Bloch-Wigner exact sequence over a semilocal domain. This exact sequence studies the third homology of $\mathit{{\rm E}}_{2}(A)$. More precisely, let $A$ be a semilocal ring such that for any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$, the residue field $A/\mathfrak{m}$ either is infinite or when $|A/\mathfrak{m}|=p^{d}$, then $(p-1)d>6$. If $-1\in{A^{\times}}^{2}$ or $|{A^{\times}}/({A^{\times}})^{2}|\leq 4$, then we obtain the refined Bloch-Wigner exact sequence
@@
where $\mathcal{RB}(A)$ is the refined Bloch group of $A$. Here the map $\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)\rightarrow H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ is obtained from the product map
@@
which is induced by the map $\mu_{2}(A)\times\mathit{{\rm E}}_{2}(A)\rightarrow\mathit{{\rm E}}_{2}(A)$, $(-1,X)\mapsto-X$. This generalizes the main result of [17]. In [19, Question 0.1] we raised the question that if over local rings the product map
@@
is trivial, provided that $|A/\mathfrak{m}_{A}|\neq 2$?
Our main tool for handling the homology groupg of $\mathit{{\rm E}}_{2}(A)$ is a first quadrant spectral sequence obtained from the complex $Y_{\bullet}(A^{2})$. All our main results are obtained from a careful analysis of this spectral sequence.
Here we outline the organization of the present paper. In Section 1 we give a short overview of the groups $\mathit{{\rm E}}_{2}(A)\subseteq\mathit{{\rm SL}}_{2}(A)$, ${\rm GE}_{2}(A)\subseteq\mathit{{\rm GL}}_{2}(A)$, the rank one Steinberg group $\mathrm{{St}}(2,A)$ and the rank one $K_{2}$-group $\mathrm{K}_{2}(2,A)$. In Section 2, we introduce and study a $\mathit{{\rm SL}}_{2}(A)$-complex $L_{\bullet}(A^{2})$ and its $\mathit{{\rm E}}_{2}(A)$-subcomplex $Y_{\bullet}(A^{2})$. In particular we will study the $H_{0}$ and $H_{1}$ of these complexes. In Section 3, we will introduce and study our main spectral sequence. In Section 4 we study the first homology group of $\mathit{{\rm E}}_{2}(A)$ and prove the first exact sequence discussed in this introduction. In Section 5 we study the natural map $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)$. In Section 6 we will prove the refined Bloch-Wigner exact sequence which studies the third homology of $\mathit{{\rm E}}_{2}(A)$ over semilocal rings.
Notations. In this paper all rings are commutative, except possibly group rings, and have the unit element $1$. For a commutative ring $A$, let $\mathcal{G}_{A}:={A^{\times}}/({A^{\times}})^{2}$ and set
@@
The element of $\mathcal{G}_{A}$ represented by $a\in{A^{\times}}$ is denoted by $\langle a\rangle$ and set $\langle\!\langle a\rangle\!\rangle:=\langle a\rangle-1\in\mathbb{Z}[\mathcal{G}_{A}]$. For $n\in\mathbb{N}$, let $\mu_{n}(A)$ denotes the group of $n$-roots of unity in $A$. Moreover, let
@@
If $\mathcal{B}\rightarrow\mathcal{A}$ is a homomorphism of abelian groups, by $\mathcal{A}/\mathcal{B}$ we mean ${\rm coker}(\mathcal{B}\rightarrow\mathcal{A})$.
@@ 1. Elementary and special linear groups of degree two
Semilocal rings and euclidean domains are examples of ${\rm GE}_{2}$-rings [21, p. 245], [3, §2]. If $F$ is a field, then $F[X_{1},\dots,X_{d}]$, $d>1$, is not a ${\rm GE}_{2}$-ring [3, page 26].
@@
(ii) If $41\leq d<$ 4, then $\mathit{{\rm E}}_{2}(A)=\mathit{{\rm SL}}_{2}(A)$, i.e. $A$ is a ${\rm GE}_{2}$-ring.
@@
See [20, Theorem 1.5]. ∎
If $41\leq d<$ 4, then $A$ is one of the rings $\mathbb{Z}[\sqrt{-1}]$, $\mathbb{Z}[\sqrt{-2}]$, $\mathbb{Z}[\sqrt{-3}]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{-3})]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{-7})]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{-11})]$. Thus these rings are ${\rm GE}_{2}$-rings.
@@ 2. The complex of unimodular vectors
[…]
@@
[…]
$H_{0}(L_{\bullet}(A^{2}))\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}\mathbb{Z}.$
@@
For the first isomorphism see the proof of [12, Theorem 3.3]. The second isomorphism follows from the first isomorphism and Corollary 2.2. ∎
@@
(ii) If $A$ is an order in a number field which is not an imaginary quadratic ring, then $\mathit{{\rm E}}_{2}(A)$ is a normal and finite-index subgroup of $\mathit{{\rm SL}}_{2}(A)$ (see Theorem 1.2). Thus by Theorem 2.3,
$H_{0}(L_{\bullet}(A^{2}))\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}\mathbb{Z}\simeq\mathbb{Z}[\mathit{{\rm SL}}_{2}(A)/\mathit{{\rm E}}_{2}(A)].$
@@
The first isomorphism follows from [12, Theorem 7.2]. The second isomorphism follows from the first isomorphism and Proposition 2.2. ∎
@@
Let $G$ be a group and let $C_{\bullet}$ be a complex of left $G$-modules: $C_{\bullet}:\ \ \cdots\rightarrow C_{1}\rightarrow C_{0}\rightarrow 0$. The $n$-th homology of $G$ with coefficient in $C_{\bullet}$, denoted by
@@ 3. The main spectral sequence and the homology of $\mathit{{\rm E}}_{2}$
This follows from an easy analysis of the spectral sequence $\mathsf{E}_{p,q}^{2}(G)$. ∎
Thus if $C_{\bullet}$ is exact in any $i\geq 1$, then we have the spectral sequence
(3.1) $E_{p,q}^{1}(G)=H_{q}(G,C_{p})\Rightarrow H_{p+q}(G,H_{0}(C_{\bullet})).$
@@
[…]
@@
where ${\boldsymbol{e}_{1}}:={\left(\!\!\!\begin{array}[]{cc}1\\ 0\end{array}\!\!\!\right)}$ and ${\boldsymbol{e}_{2}}:={\left(\!\!\!\begin{array}[]{cc}0\\ 1\end{array}\!\!\!\right)}$. We choose $({\boldsymbol{\infty}})$ and $({\boldsymbol{\infty}},{\boldsymbol{0}})$ as representatives of the orbit of the generators of $Y_{0}(A^{2})$ and $Y_{1}(A^{2})$, respectively. Then
@@
$E_{0,q}^{1}\simeq H_{q}(\mathrm{B}(A),\mathbb{Z}),\ \ \ \ \ \ E_{1,q}^{1}\simeq H_{q}(\mathrm{T}(A),\mathbb{Z}).$
In particular, $E_{0,0}^{1}\simeq\mathbb{Z}\simeq E_{1,0}^{1}$. Moreover, $d_{1,q}^{1}=H_{q}(\sigma)-H_{q}({\rm inc})$, where $\sigma:\mathrm{T}(A)\rightarrow\mathrm{B}(A)$ is given by $\sigma(D(a))=wD(a)w^{-1}=D(a^{-1})$ for $w={\left(\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\!\right)}$. This easily implies that $d_{1,0}^{1}$ is trivial and
@@
$E_{0,1}^{2}\simeq\mathcal{G}_{A}\oplus A_{A^{\times}},\ \ \ \ker(d_{1,1}^{1})\simeq\mu_{2}(A).$
@@
it follows that $d_{2,0}^{1}:H_{0}(\mathit{{\rm E}}_{2}(A),Z_{1}^{Y}(A^{2}))\rightarrow\mathbb{Z}$ is surjective. Hence $E_{1,0}^{2}=0$. Let
@@
We denote the kernel of $\epsilon$ by $I(A)$. Let
@@
is given by $[x]\mapsto\langle\!\langle x\rangle\!\rangle\langle\!\langle 1-x\rangle\!\rangle$. All these induce a surjective map

### v3

@@ Abstract.
In this article, we study the first, second, and third homology groups of the elementary group $\mathit{{\rm E}}_{2}(A)$, where $A$ is a commutative ring. In particular, we establish a refined Bloch–Wigner type exact sequence over a semilocal ring (subject to mild restrictions on its residue fields) in which either $1\in({A^{\times}})^{2}$ or $|{A^{\times}}/({A^{\times}})^{2}|\leq 4$.
For a commutative ring $A$ with unit, let $\mathit{{\rm E}}_{2}(A)$ denote the subgroup of the general linear group $\mathit{{\rm GL}}_{2}(A)$ generated by elementary matrices. Clearly, $\mathit{{\rm E}}_{2}(A)\subseteq\mathit{{\rm SL}}_{2}(A)$. We say that $A$ is a ${\rm GE}_{2}(A)$-ring if $\mathit{{\rm E}}_{2}(A)=\mathit{{\rm SL}}_{2}(A)$.
In two previous papers [17] and [18] we studied the low-dimensional homology groups of the elementary group $\mathit{{\rm E}}_{2}(A)$. In this paper, we further investigate these homology groups.
The first main result of this paper concerns the first homology group of $\mathit{{\rm E}}_{2}(A)$. We show that for any commutative ring $A$, there is an exact sequence
@@
where $Y_{\bullet}(A^{2})$ is a complex of ${\rm GE}_{2}(A)$-modules obtained from certain unimodular vectors in $A^{2}$ (see Theorem 4.1). Here $M$ is the additive subgroup of $A$ generated by $x(a^{2}-1)$ and $3(b+1)(c+1)$, where $x\in A$ and $a,b,c\in{A^{\times}}$. The rightmost map is induced by the homomorphism $A\rightarrow\mathit{{\rm E}}_{2}(A)$ sending $x$ to the elementary matrix $E_{12}(x)$. The group $A/M$ and its connection with the first homology of $\mathit{{\rm E}}_{2}(A)$ were first studied by Cohn in [3] and [4].
Our second main result concerns the second homology group $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$. For simplicity, let $A$ be a semilocal ring, where all its residue fields have more than three elements. Then it is not difficult to show that this homology group fits into the exact sequence
@@
where $\mathrm{B}(A)$ is the subgroup of $\mathit{{\rm E}}_{2}(A)$ consisting of all matrices whose lower left-hand entry is equal to zero. Here $I(A)$ denotes the fundamental ideal of $A$. More precisely, let $\mathcal{G}_{A}$ be the square class group of $A$ and $\mathcal{I}_{A}$ the augmentation ideal of $\mathbb{Z}[\mathcal{G}_{A}]$. Let $\langle a\rangle$ denotes the element of $\mathcal{G}_{A}$ represented by $a\in{A^{\times}}$. Then we define
$I(A):=\mathcal{I}_{A}/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a,1-a\in{A^{\times}}\rangle,$
where $\langle\!\langle a\rangle\!\rangle:=\langle a\rangle-1\in\mathbb{Z}]\mathcal{G}_{A}]$. Observe that $I^{2}(A):=(I(A))^{2}$. Our second result concerns the map $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)$ appearing in the above exact sequence. When $2\in{A^{\times}}$, we exhibit a cycle whose homology class in $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ maps to the element $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle$, for $a,b\in{A^{\times}}$ (see Proposition 5.2 for the precise statement).
The third main result of this paper is a refined Bloch–Wigner exact sequence over a semilocal domain, which describes the third homology of $\mathit{{\rm E}}_{2}(A)$. More precisely, let $A$ be a semilocal ring such that for any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$, the residue field $A/\mathfrak{m}$ is either infinite, or if $|A/\mathfrak{m}|=p^{d}$, then (p1)d>6(p-1)d>6. If $1\in{A^{\times}}^{2}$ or $|{A^{\times}}/({A^{\times}})^{2}|\leq 4$, then we obtain the refined Bloch–Wigner exact sequence
@@
where $\mathcal{RB}(A)$ is the refined Bloch group of $A$ (see Theorem 6.3). Here the map $\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)\rightarrow H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ is obtained from the product map
@@
which is induced by the map $\mu_{2}(A)\times\mathit{{\rm E}}_{2}(A)\rightarrow\mathit{{\rm E}}_{2}(A)$, $(-1,X)\mapsto-X$. This generalizes the main result of [18]. In [20, Question 0.1], we raised the question of whether, over local rings, the product map
@@
is trivial, provided that $|A/\mathfrak{m}_{A}|\neq 2$.
Our main tool for studying the homology groups of $\mathit{{\rm E}}_{2}(A)$ is a first-quadrant spectral sequence derived from the complex $Y_{\bullet}(A^{2})$. All our main results are obtained from a detailed analysis of this spectral sequence.
We now outline the organization of the present paper. In Section 1 we give a brief overview of the groups $\mathit{{\rm E}}_{2}(A)\subseteq\mathit{{\rm SL}}_{2}(A)$, ${\rm GE}_{2}(A)\subseteq\mathit{{\rm GL}}_{2}(A)$, the rank-one Steinberg group $\mathrm{{St}}(2,A)$, and the rank-one $\mathrm{K}_{2}$-group $\mathrm{K}_{2}(2,A)$. In Section 2, we introduce and study an $\mathit{{\rm SL}}_{2}(A)$-complex $L_{\bullet}(A^{2})$ and its $\mathit{{\rm E}}_{2}(A)$-subcomplex $Y_{\bullet}(A^{2})$. In particular, we study the groups $H_{0}$ and $H_{1}$ of these complexes. In Section 3, we introduce and analyze our main spectral sequence. In Section 4, we study the first homology group of $\mathit{{\rm E}}_{2}(A)$ and prove the first exact sequence discussed in this introduction. In Section 5, we study the natural map $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)$. In Section 6, we establish the refined Bloch–Wigner exact sequence, which describes the third homology of $\mathit{{\rm E}}_{2}(A)$ over semilocal rings.
Notations. Throughout this paper all rings are commutative, except possibly group rings, and possess a unit element $1$. For a commutative ring $A$, let $\mathcal{G}_{A}:={A^{\times}}/({A^{\times}})^{2}$ and define
@@
The element of $\mathcal{G}_{A}$ represented by $a\in{A^{\times}}$ is denoted by $\langle a\rangle$, and we set $\langle\!\langle a\rangle\!\rangle:=\langle a\rangle-1\in\mathbb{Z}[\mathcal{G}_{A}]$. For $n\in\mathbb{N}$, let $\mu_{n}(A)$ denote the group of $n$th roots of unity in $A$. Moreover, define
@@
If $\mathcal{B}\rightarrow\mathcal{A}$ is a homomorphism of abelian groups, we write $\mathcal{A}/\mathcal{B}$ for ${\rm coker}(\mathcal{B}\rightarrow\mathcal{A})$.
@@ 1. Elementary and special linear groups of degree two
Semilocal rings and euclidean domains are examples of ${\rm GE}_{2}$-rings [23, p. 245], [3, §2]. If $F$ is a field, then $F[X_{1},\dots,X_{d}]$, $d>1$, is not a ${\rm GE}_{2}$-ring [3, page 26].
@@
(ii) If $1\leq d<4$, then $\mathit{{\rm E}}_{2}(A)=\mathit{{\rm SL}}_{2}(A)$, i.e. $A$ is a ${\rm GE}_{2}$-ring.
@@
See [22, Theorem 1.5]. ∎
If $1\leq d<4$, then $A$ is one of the rings $\mathbb{Z}[\sqrt{-1}]$, $\mathbb{Z}[\sqrt{-2}]$, $\mathbb{Z}[\sqrt{-3}]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{-3})]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{-7})]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{-11})]$. Thus these rings are ${\rm GE}_{2}$-rings.
@@ 2. The complex of unimodular vectors
Let $\partial_{-1}=\varepsilon:L_{0}(A^{2})\rightarrow\mathbb{Z}$ be the natural augmentation map, which is given by $\sum_{i}n_{i}(\langle{\boldsymbol{v}_{0,i}}\rangle)\mapsto\sum_{i}n_{i}$. Then we have the complexes $L_{\bullet}(A^{2})\overset{\varepsilon}{\rightarrow}\mathbb{Z}$ and $Y_{\bullet}(A^{2})\overset{\varepsilon}{\rightarrow}\mathbb{Z}$.
@@
For any commutative ring $A$, $H_{0}(Y_{\bullet}(A^{2}))\overset{\bar{\varepsilon}}{\simeq}\mathbb{Z}$, where $\bar{\varepsilon}$ is induced by $\varepsilon$. In other words, the complex $Y_{1}(A^{2})\overset{\partial_{1}^{Y}}{\longrightarrow}Y_{0}(A^{2})\overset{\varepsilon}{\rightarrow}\mathbb{Z}\rightarrow 0$ always is exact. Moreover,
$H_{0}(L_{\bullet}(A^{2}))\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}\mathbb{Z}\simeq\mathbb{Z}[\mathit{{\rm SL}}_{2}(A)/\mathit{{\rm E}}_{2}(A)],$
where the right side is the coset module.
@@
For the first isomorphism see the proof of [12, Theorem 3.3]. The isomorphism $H_{0}(L_{\bullet}(A^{2}))\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}\mathbb{Z}$ follows from the first isomorphism and Corollary 2.2. The other isomorphism follows from the fact that $\mathbb{Z}[\mathit{{\rm SL}}_{2}(A)]$ is a free $\mathbb{Z}[\mathit{{\rm E}}_{2}(A)]$-module with a basis consisting of one representative for each element of the set $\mathit{{\rm SL}}_{2}(A)/\mathit{{\rm E}}_{2}(A)$.
@@
(ii) If $A$ is an order in a number field which is not an imaginary quadratic ring, then $\mathit{{\rm E}}_{2}(A)$ is a normal and finite-index subgroup of $\mathit{{\rm SL}}_{2}(A)$ (see Theorem 1.2). Hence $H_{0}(L_{\bullet}(A^{2}))$ is a free abelian group of finite rank.
@@
The first isomorphism follows from [12, Theorem 7.2]. The second isomorphism follows from the first isomorphism and Proposition $2.$ ∎
@@
Let $G$ be a group and let $C_{\bullet}$ be a complex of left $G$-modules: $C_{\bullet}:\ \ \cdots\rightarrow C_{1}\rightarrow C_{0}\rightarrow 0$. The $n$-th homology of $G$ with coefficients in $C_{\bullet}$, denoted by
@@ 3. The main spectral sequence and the homology of $\mathit{{\rm E}}_{2}$
Apply [2, Chap. VII, Proposition 5.2] to the morphism of complexes $C_{\bullet}\rightarrow M_{\bullet}$, where $M_{\bullet}$ is the complex consisting of the single module $M$ concentrated in degree 0. ∎
Thus if $C_{\bullet}$ is exact for every $i\geq 1$, then we have the spectral sequence
(3.1) $\mathrm{E}_{p,q}^{1}(G)=H_{q}(G,C_{p})\Rightarrow H_{p+q}(G,H_{0}(C_{\bullet})).$
@@
$0\rightarrow Z_{1}^{Y}(A^{2})\overset{{\rm inc}}{\rightarrow}Y_{1}(A^{2})\rightarrow Y_{0}(A^{2})\overset{\varepsilon}{\rightarrow}\mathbb{Z}\rightarrow 0,$
@@
${\boldsymbol{\infty}}:=\langle{\boldsymbol{e}_{1}}\rangle,\ \ \ {\boldsymbol{0}}:=\langle{\boldsymbol{e}_{2}}\rangle,\ \ \ {\boldsymbol{a}}:=\langle{\boldsymbol{e}_{1}}+a{\boldsymbol{e}_{2}}\rangle,\ \ \ a\in{A^{\times}},$
@@
$\mathrm{E}_{0,q}^{1}\simeq H_{q}(\mathrm{B}(A),\mathbb{Z}),\ \ \ \ \ \ \mathrm{E}_{1,q}^{1}\simeq H_{q}(\mathrm{T}(A),\mathbb{Z}).$
In particular, $\mathrm{E}_{0,0}^{1}\simeq\mathbb{Z}\simeq\mathrm{E}_{1,0}^{1}$. Moreover, $d_{1,q}^{1}=H_{q}(\sigma)-H_{q}({\rm inc})$, where $\sigma:\mathrm{T}(A)\rightarrow\mathrm{B}(A)$ is given by $\sigma(D(a))=wD(a)w^{-1}=D(a^{-1})$ for $w={\left(\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\!\right)}$. This easily implies that $d_{1,0}^{1}$ is trivial and
@@
$\mathrm{E}_{0,1}^{2}\simeq\mathcal{G}_{A}\oplus A_{A^{\times}},\ \ \ \ker(d_{1,1}^{1})\simeq\mu_{2}(A).$
@@
it follows that $d_{2,0}^{1}:H_{0}(\mathit{{\rm E}}_{2}(A),Z_{1}^{Y}(A^{2}))\rightarrow\mathbb{Z}$ is surjective. Hence $\mathrm{E}_{1,0}^{2}=0$. Let
@@
We denote the kernel of $\epsilon$ by $I(A)$.
Combining all these results, we obtain that the first page of the spectral sequence has the following form:
[…]
The second page of the spectral sequence is given by:
[…]
Let
@@
is given by $x]\mapsto\langle\!\langle x\rangle\!\rangle\langle\!\langle 1-x\rangle\!\rangle$. All these induce a surjective map